# KaFT: Knowledge-aware Fine-tuning for Boosting LLMs’ Domain-specific Question-Answering Performance

### v1

@@ Abstract
Qihuang Zhong1, Liang Ding2, Xiantao Cai1, Juhua Liu1thanks:   Corresponding Authors: Juhua Liu (e-mail: liujuhua@whu.edu.cn), Bo Du (e-mail: dubo@whu.edu.cn), Bo Du1∗, Dacheng Tao3 1 School of Computer Science, National Engineering Research Center for Multimedia Software, Institute of Artificial Intelligence and Hubei Key Laboratory of Multimedia and Network Communication Engineering, Wuhan University, China 2 The University of Sydney, Australia  3 Nanyang Technological University, Singapore {zhongqihuang, caixiantao, liujuhua, dubo}@whu.edu.cn {liangding.liam, dacheng.tao}@gmail.com
@@ 2.3 Empirical Analyses
Table 1: Fine-grained test results of trained LLaMA3-8B models on MedQA. The first column denotes the models trained with different subsets, and the first row denotes the different test subsets. “mig-ri” and “mig-wr” refer to might-right and might-wrong subsets, respectively.
@@
Intuitively, a straightforward way for alleviating the negative effect of knowledge conflict is to filter out the wrong subset from $\mathcal{D}$ and use the other subsets for SFT. However, as aforementioned, appropriate conflict data might help activate LLMs’ internal knowledge and lead to better performance. To verify our conjecture, we introduce different proportions $\lambda$ of conflict data from wrong into the collection of other less-conflict subsets, where $\lambda$ ranges from 0% to 100%. The results are illustrated in Figure 2 (c), from which we find that compared to directly filtering the wrong (i.e., $\lambda=0\%$), introducing some conflict data (e.g., $\lambda=50\%$) could be more beneficial. This highlights the necessity of exploring more effective SFT methods to make full use of the conflict data.
@@ 4.2 Compared Results
Additionally, we evaluate the tuned LLMs on the OOD benchmarks to verify LLMs’ robustness. Due to space limitations, we only present the contrastive results of Mistral-7B and LLaMA3-3B models in Table 3. From the table, we observe that KaFT significantly outperforms the baselines among all domains, indicating that alleviating the negative effect of conflict data can avoid the overfitting of LLMs, continuing to prove the effectiveness of our proposed KaFT approach.
@@ 6 Related Works
Different from these prior studies, we first design a query diversification strategy to robustly detect the conflict and then propose the KaFT to make full use of all training data. The main idea of KaFT is to use sample-adaptive rewards for better guiding the learning of LLMs, which is somewhat similar to prior adaptive-learning methods Wang et al. (2024a); Li et al. (2020); Kang et al. (2020); Ghorbani and Zou (2019); Yoon et al. (2020); Zhong et al. (2024). We should emphasize that, although assigning different weights to subsets is a common and intuitive approach, it is non-trivial to determine the weights for different subsets, especially for the domain-specific SFT of LLMs. The way of KaFT to obtain the sample-adaptive rewards is innovative, and thus we believe that our work is insightful and has great potential to unleash the power of LLMs in real-world applications.
@@ A.1 Details of Tasks and Datasets
A.2 Prompt Details

### v2

@@ Abstract
Qihuang Zhong1, Liang Ding2, Xiantao Cai1∗, Juhua Liu1thanks:   Corresponding Authors: Xiantao Cai (e-mail: caixiantao@whu.edu.cn), Juhua Liu (e-mail: liujuhua@whu.edu.cn), Bo Du1, Dacheng Tao3 1 School of Computer Science, National Engineering Research Center for Multimedia Software, Institute of Artificial Intelligence and Hubei Key Laboratory of Multimedia and Network Communication Engineering, Wuhan University, China 2 The University of Sydney, Australia  3 Nanyang Technological University, Singapore {zhongqihuang, caixiantao, liujuhua, dubo}@whu.edu.cn {liangding.liam, dacheng.tao}@gmail.com
@@ 2.3 Empirical Analyses
Table 1: Fine-grained test results of trained LLaMA3-8B models on MedQA. “mig-ri” and “mig-wr” refer to might-right and might-wrong subsets.
@@
Intuitively, a straightforward way for alleviating the negative effect of knowledge conflict is to filter out the wrong subset from $\mathcal{D}$ and use the other subsets for SFT. However, as aforementioned, appropriate conflict data might help activate LLMs’ internal knowledge and lead to better performance. To verify our conjecture, we introduce different proportions $\lambda$ of conflict data from wrong into the collection of other less-conflict subsets, where $\lambda$ ranges from 0% to 100%. The results are illustrated in Figure 2 (c), from which we find that compared to directly filtering the wrong (i.e., $\lambda=0\%$), introducing some conflict data (e.g., $\lambda=25\%$) could be more beneficial. This highlights the necessity of exploring more effective SFT methods to make full use of the conflict data.
@@ 4.2 Compared Results
Additionally, we evaluate the tuned LLMs on the OOD benchmarks to verify LLMs’ robustness. Due to space limitations, we only present the contrastive results of Mistral-7B and LLaMA3-3B models in Table 3. It can be observed that KaFT significantly outperforms the baselines among all domains, indicating that alleviating the negative effect of conflict data can avoid the overfitting of LLMs, continuing to prove KaFT’s effectiveness.
@@ 6 Related Works
Different from these prior studies, we first design a query diversification strategy to robustly detect the conflict and then propose KaFT to make full use of all training data. The main idea of KaFT is to use sample-adaptive rewards for better guiding the learning of LLMs, which is somewhat similar to prior adaptive-learning methods Wang et al. (2024a); Li et al. (2020); Kang et al. (2020); Ghorbani and Zou (2019); Yoon et al. (2020); Zhong et al. (2024). We should emphasize that, although assigning different weights to subsets is a common and intuitive approach, it is non-trivial to determine the weights for different subsets, especially for the domain-specific SFT of LLMs. Thus, we believe that KaFT is innovative and our work is insightful.
@@ 7 Conclusion
Acknowledgements
We are grateful to the anonymous reviewers and the area chair for their insightful comments and suggestions. This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2705700, in part by the National Natural Science Foundation of China under Grant 623B2076, U23B2048 and 62225113, in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017, and in part by the Science and Technology Major Project of Hubei Province under Grant 2024BAB046. Dr Tao’s research is partially supported by NTU RSR and Start Up Grants. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
@@ A.1 Details of Tasks and Datasets
A.2 Training and Evaluation Details
A.3 Prompt Details